OpenStudy (anonymous):

a^5=b^6,c^3=d^4, and d-a=61,find the value of c-b

OpenStudy (shubhamsrg):

all are natural integers ? is it given ?

OpenStudy (shubhamsrg):

natural nos *

OpenStudy (anonymous):

positive intergers

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

using your formula for a and d (jonas0 if we let x = (b/c and y = (a.b) we get (x - y^2)(x^2 + xy^2 + y^4) -61 = 0 solving this might help

OpenStudy (anonymous):

is it wrong to say \[64-3=61\] d=64,a=3

OpenStudy (cwrw238):

is that just a guess?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I mean positive integer.

OpenStudy (anonymous):

so can we guess until b is positive int

OpenStudy (cwrw238):

@jonas - you deleted those values for a and b - have i got them right

OpenStudy (anonymous):

I guess yes.... But guessing is not a nice method.

OpenStudy (anonymous):

\[a=(\frac{ a }{ b })^6,d=(\frac{ c }{ d })^3\]

OpenStudy (cwrw238):

solution of that equation in x and y is x = 5 , y = -2, or 2

OpenStudy (cwrw238):

a/b = 5, c/d = 2 or -2

OpenStudy (cwrw238):

a = 5b, c = 2d or c = -2d - not sure how to go forward from here....

OpenStudy (anonymous):

Did u get 593 @cwrw238

OpenStudy (cwrw238):

trial and error would be tedious i guess

OpenStudy (cwrw238):

593?

OpenStudy (cwrw238):

oh for c - b no i've only got as far as the last relations i posted

OpenStudy (anonymous):